It considers that the following equations represent the supply and demand of honey in the national market. Answer the following questions:

?? = 400−2? ?? = 1/2?

a. Find the market balance.

b. Calculate the consumer and producer surplus.
Answer #1

a. Market balance /equilibrium occurs when the quantity demanded equals the quantity supplied

i.e Qd=Qs

Substituting in the above equation

400-2P=1/2P

800-4P=P

800=5P

P=160

Substituting the value of P in Qd,we get

=400-2(160)

=400-320

=80

Therefore the equilibrium price of honey is 160 and the equilibrium quantity of honey demanded is 80

b. Consumer surplus is the difference between the total amount that a consumer is willing and able to pay for a good or service and the amount that they actually pay for the good or service .

Consumer surplus=(1/2)*80*40=1600

Producer surplus is the difference between the amount that a producer is willing and able to supply a good for and the price that they actually receive

Producer surplus=(1/2)*80*160=6400